A 10kW, 400 V, 4pole, delta connected SCIM gave the following test results,

N.L. test: 250W, 400V, 8A, B.R. Test : 1350W, 90V, 35A ,

The stator winding resistance/phase is 0.72 ohms/phase. Calculate rotational losses and the

equivalent circuit parameters of the motor.


1
Expert's answer
2021-06-07T05:46:20-0400

"Reff = 1.2*r;"  Due to delta connected rotor. "Reff=1.2*0.6=0.72 \Omega"

"Pr=250\u22123\u2217 \n3\n64\n\u200b\n \u2217(0.6\u22171.2)=204W."